# AcousTools: A ‘Full-Stack’, Python-Based, Acoustic Holography Library

### v3

@@ II.2 Existing Software
Propagators Piston Model GSPAT Y Y Y Y
Propagators BEM HighSpeedBEM Y N N N
Solvers Naive VolDisplaySingleParticle Y Y Y N
Solvers IBIB Y Y N N
Solvers GS-PATGSPAT Y Y N N
Solvers Gradient DescentAcousticElements Y N N Y
Solvers Implicit-Gradient Diff-PAT Y N N N
@@
Analysis Gor’kov Potential AcousticElements Y Y Y Y
Analysis Force AcousticElements Y Y Y Y
Analysis Stiffness AcousticElements Y Y N Y
@@
Table 1: Summary of various acoustic holography software library and packages. AcousTools meets each step of the full-stack discussed above while each other solution fails at at least part of some of the steps. AcousTools is unique in that it can be used as a complete solution that can swap and change between different propagators, solvers and contexts for various applications.
@@ II.2.1 Ultraino
Ultraino is an early Java-based software package that aimed to allow for start to finish acoustic development, being able to generate phases, drive transducer arrays, and simulate the resulting sound field Ultraino. This package allows for the simulation and control of sound fields in free-field scenarios due to its simple propagation model and naive focusing method. Their framework allows for many different transducer arrays to be used, even with more obscure geometries such as a convex curved array. It also supports the computation compute some metrics, such as pressure, Gor’kov potential, and force, while lacking a stiffness measure.
@@
However, Ultraino is a stand-alone application compared to a code library, and it is not clear how a user would extend the software to support new use cases or integrate it with existing code, as no instructions are provided for doing so UltrainoGitHub. Furthermore, there is limited support for more advanced methods used in acoustic holography, such as complex multipoint algorithms like GS-PAT or propagators that allow scattering computation like BEM.
II.2.2 OpenMPD
@@
OpenMPD was a later solution aiming to be a “low-level presentation engine” focusing on a high-speed rendering pipeline and giving access to a Unity interface for ease of use OpenMPD. This engine supports 10kHz update rates, sufficient for persistence of vision effectsGSPAT, and is able to run lower frequency haptic content, giving them a wide variety of utility. It also provides an OpenCL implementation of the fast solver GS-PATGSPAT as well as code to compute all four analysis quantities that have been discussed above, from pressure to stiffness.
@@
OpenMPD does not allow for arbitrary arrays by default, being limited to flat 16x16 arrays. While the developers note in their paper that different configurations could be implemented, this would require changes to the OpenMPD code. As such, this does not constitute native support for arbitrary geometries. Additionally, only the piston model propagator is supplied, and thus the more complex scenarios that BEM unlocks HapticSoftness; HighSpeedBEM are not possible with OpenMPD as provided. Similarly, only the projective solvers (i.e., GS-PAT, naive and IB) are provided, meaning any scenario that requires a different solver cannot be used.
II.2.3 Levitate
@@
Levitate is a python-based acoustic holography library Levitate, which allows modelling propagation from transducer arrays to points in space based on the piston model. This library supports an explicit-gradient descent solver where the gradient must be computed by the user and does not integrate any of the more common projective methods or allow for implicit-gradient methods, which may limit its application as these are the dominant methods used across various domains HapticsMPSTM; LeviPrint; Articulev. Levitate can compute all of the analysis metrics discussed but has no implementation for hardware support, indicating that it is strictly a software only library. Finally, it also lacks BEM support, giving it similar disadvantages discussed above. This could be a useful software package and maybe one of the most comparable to AcousTools, however, the lack of support for critical sections of the modern holography pipeline, such as many of the most common solvers, would limit its utility.
II.2.4 Other Software
@@
There are various other pieces of acoustic software, which mostly aim to simulate acoustic wave propagation and lack most of the features that we have identified that a developer would need. For example, K-wave is a MATLAB package that enables simulation of free-field acoustics and definition of arbitrary arrays. However, it only supports the most simple geometric solvers, does not have any scattering support and cannot compute any of the more complex analysis metricsKwave. Similar issues are found in other software implementations, such as Ultrasim ultrasim and FOCUS FOCUS, which all lack the versatility that would be needed to give a developer the full range of options.
There are other packages that aim to give BEM implementations for modelling ultrasound scattering such as OpenBEMOpenBEM which implements BEM but has no support for advanced solvers, analysis, lacks documentation and has no public download. In a similar vein BEM++BEMpp lacks these same features but does have documentation. As well as these issues, none of these packages give hardware support - limiting them to simulation only uses.
@@
Because of the aforementioned reasons, these packages are not considered in the remaining of this work as they do not provide the user with enough of the full-stack to allow for flexible development. For a (non-exhastive) list of acoustic simulation software, which share many of these issues, see Ref.kwave-software-list.
@@
Alternatively, some software acts as a single application for acoustic development instead of a library for development. One such pipeline is ArticuLevArticulev, which allows developers to use acoustically levitated props to create display content. ArticuLev provides support for the creation of traps and props and it’s methods are agnostic to the underlying algorithms that are used to generate these traps but only a multipoint version of the naive algorithm is actually provided and in a similar way, they are agnostic to the specific array but assume a 16x16 array. This is not strictly comparable to AcousTools as this is a platform built for a specific application instead of a general purpose software library but should be mentioned that for developers with specific goals, specific software may exist Articulev.
III The Acoustic ‘Full-Stack’
@@ III.0.1 Setup
The bottom of the stack is the setup phase, defining the transducer array’s shape and parameters as well as the locations in space where acoustic waves are focused LeviPrint; AcousticElements.
@@
The transducer arrays used across acoustic applications vary from flat square boards OpenMPD to rectangle AUTD3 and curved surfaces TinyLev. Even more complex geometries have also been investigated in theory, such as spherical or triangular arrays, but these are rarely used in practice due to the added complexity LeviPrint. Because of the variety of different options, the first decision that a developer must make is which of these arrays to use, and the software they use should accommodate this.
@@
After defining the array, still as part of setup, the developer needs to conceptually define where in space the acoustic waves should be focused. The area a given transducer array can focus will vary based on its geometry. For example, a single-axis device TinyLev has significantly less freedom in focusing locations compared to a larger 3D array OpenMPD; AUTD3. Therefore, the transducer geometry needs to be accounted for when defining where the focal points can be created.
@@
Finally, the various acoustic parameters and constants need to be defined, such as the frequency of sound (most often 40kHzOpenMPD; Ultraino; AUTD3; GSPAT; VolDisplaySingleParticle; HapticsMPSTM; LeviPrint), the reference pressure of transducers at 1m GSPAT, the density of the medium and levitated objects, and the speed of sound in that medium BruusAcoustofluidics. These are often not choices a developer would actively make, but rather physical properties of the system they have selected, for example the frequency of waves used are dependant on the resonant frequency of the transducers. However, these parameters should be acknowledged, as earlier design decisions will influence them.
III.0.2 Propagators
@@
Once the transducers and points have been defined, the next step is to couple them together by modelling how acoustic waves travel from transducers to points GSPAT; Diff-PAT; HapticsMPSTM; HighSpeedBEM; ToBEMorNot; HapticSoftness. There are two main scenarios that can be used, free-field with no scattering objects and propagation with obstacles acting as scatterers. The free-field case, using the piston model (Eq.1), is fairly easy to compute and can be calculated at very high speeds for real-time dynamic holography GSPAT. It is versatile and, in most free-field cases, the scattering effects can be neglected as they do not impact performanceGSPAT; LeviPrint.
@@
However, in some cases, scattering effects are too significant to ignore HighSpeedBEM; ToBEMorNot. Thus, the developer may choose to use a propagator that models this scattering. Given the model of the scatterer as a mesh, BEM can be used to predict how the sound scatterers off the boundary of the object HighSpeedBEM as seen in Eq.3. Accounting for them can bring improvements in a multitude of applications HapticSoftness; ToBEMorNot; HighSpeedBEM but come at a significant computational cost ToBEMorNot.
@@ III.0.3 Solvers
Once the setup and propagator have been defined, acoustic solvers compute the transducer activations that cause transducers to focus sound at the chosen points VolDisplaySingleParticle; GSPAT; ToBEMorNot. Supporting a number of solvers enables a developer to have different options, but determining the best solver requires careful analysis of the problem the developer is trying to solve. For an application that requires fast computation speeds, an algorithm such as GS-PATGSPAT may be best. On the other hand, WGS TGS gives high pressure and low variance across multiple points. For more complex scenarios, a implicit-gradient descent optimiser can solve more complex objectives Diff-PAT. It is important for a developer to be able to define custom objective functions for their own context and the developer must weight which is best, potentially comparing multiple options before a single best choice is decided upon. This emphasises the desirability for simple swapping between solvers, to aid in this development process.
III.0.4 Analysis
@@
At this point, a developer has defined their setup, propagators, and solvers to generate acoustic focal points at the desired locations. After this, they may need to quantify certain properties of the sound field (e.g., pressure, Gor’kov potential, force, stiffness), either for validation or further optimisation AcousticElements; AcoustoFab; HighSpeedBEM; LeviPrint. Importantly, this step may overlap with the Solvers step. For example, using a custom objective defined to maximise the pressure at a point may require elements from both the previous and current layer. However, this layer can also sit by itself, for example, quantifying the force on a particle to validate if a particle can be correctly levitated at the desired location Ultraino. The quantities discussed here, which help to quantify an acoustic field, are pressure as well as Gor’kov potential, force, and stiffness as described in Eqs. 8, 11, and 12 respectively.
@@
Which one of these metrics is the most useful will depend on the use case a specific developer might have. If they are developing a haptic interface, the pressure at a point might be the only thing that they need to consider HapticsMPSTM while for levitation the Gor’kov potential, force, and stiffness might become more appropriate HighSpeedBEM; AcousticElements. Additionally, the square of pressure correlates very well with stiffness, and as such, the simpler metric can be used as a proxy for the more complex one in practice GSPAT. As discussed above, having a diverse set of tools means that the correct tool can be used at the right moment.
III.0.5 Hardware
@@
The final step before an acoustic holography application can be deployed is to control the physical transducers in the real world. This involves sending synchronised phase and amplitude values from a computer to the device, physically realising the until-this-point simulated fields. This presents difficulties as for the very high update rates that acoustic holography often demands, up to 10kHz for levitation, a high degree of synchronization is needed. Various platforms exist to address this problem. Such platforms include OpenMPDOpenMPD, which can run at 10kHz; AUTD3 AUTD3, which is able to run at 1kHz; and UltrainoUltraino, which runs at 25kHz. These different platforms also have different ways to communicate between the transducer arrays and the computer, which would need to be accommodated by any software used.
@@
This section is more difficult for software packages to be general purpose, as each different hardware setup has a different communication method and structure for the messages to be sent OpenMPD; AUTD3. Thus, many software packages only function for one or a small number of hardware devices but having at least one supported hardware device enables the simulated fields to be physicalised in real world.
IV AcousTools
@@
As discussed above, the existing software implementations for acoustic holography are limited in their functionality, often not supporting a full range of solvers or propagators. To address this, we present AcousTools; a python library built to meet the full-stack of acoustic holography. AcousTools is open source and can be found at our GitHub repositoryAcousTools.
@@
Python was chosen to implement AcousTools because of its reputation for accessibility, especially for novice users pythonEase. It is one of the most widely used programming languages PythonUse and among the fastest for beginners to learn. Whereas C++, commonly used in acoustic holography, is considered one of the most difficult PythonLearning. This popularity, combined with Python’s ease of adoption and of use, suggests that adoption and extension of AcousTools would be similarly straightforward.
@@
Additionally, AcousTools is based on PyTorchPyTorch due to its parallelised Tensor implementations, GPU support, and automatic differentiation system. These together make PyTorch an ideal base for AcousTools and provide the secondary benefit that potential future machine leaning applications for acoustic holography could be easily implemented, as PyTorch is widely used in this domain.
@@
Here we discuss AcousTools’ ability to meet every step of every stage of the acoustic full-stack discussed above. Table 1 summarises both the existing software’s performance and AcousTools’ features, illustrating that AcousTools meet every stage while the others all lack in some area or another. There is also a lack of standardisation across acoustic holography research OpenMPD; AUTD3; Ultraino, and we believe that AcousTools with its general purpose nature could fill this gap for researchers across all domains.
IV.1 Setup

### v6

@@ 2.2 Existing Software
Propagators Piston Model [GSPAT] Y Y Y Y
Propagators BEM [HighSpeedBEM] Y N N N
Solvers Naive [VolDisplaySingleParticle] Y Y Y N
Solvers IB[IB] Y Y N N
Solvers GS-PAT[GSPAT] Y Y N N
Solvers Gradient Descent[AcousticElements] Y N N Y
Solvers Implicit-Gradient [Diff-PAT] Y N N N
@@
Analysis Gor’kov Potential [AcousticElements] Y Y Y Y
Analysis Force [AcousticElements] Y Y Y Y
Analysis Stiffness [AcousticElements] Y Y N Y
@@
Table 1: Summary of various acoustic holography software libraries and packages. AcousTools meets each step of the full-stack discussed above while each other solution fails at at least part of some of the steps. AcousTools is unique in that it can be used as a complete solution that can swap and change between different propagators, solvers and contexts for various applications.
@@ 2.2.1 Ultraino
Ultraino is an early Java-based software package that aimed to allow for start to finish acoustic development, being able to generate phases, drive transducer arrays, and simulate the resulting sound field [Ultraino]. This package allows for the simulation and control of sound fields in free-field scenarios due to its simple propagation model and naive focusing method. Their framework allows for many different transducer arrays to be used, even with more obscure geometries such as a convex curved array. It also supports the computation compute some metrics, such as pressure, Gor’kov potential, and force, while lacking a stiffness measure.
@@
However, Ultraino is a stand-alone application compared to a code library, and it is not clear how a user would extend the software to support new use cases or integrate it with existing code, as no instructions are provided for doing so [UltrainoGitHub]. Furthermore, there is limited support for more advanced methods used in acoustic holography, such as complex multipoint algorithms like GS-PAT or propagators that allow scattering computation like BEM.
2.2.2 OpenMPD
@@
OpenMPD was a later solution aiming to be a “low-level presentation engine” focusing on a high-speed rendering pipeline and giving access to a Unity interface for ease of use [OpenMPD]. This engine supports 10kHz update rates, sufficient for persistence of vision effects[GSPAT], and is able to run lower frequency haptic content, giving them a wide variety of utility. It also provides an OpenCL implementation of the fast solver GS-PAT[GSPAT] as well as code to compute all four analysis quantities that have been discussed above, from pressure to stiffness.
@@
OpenMPD does not allow for arbitrary arrays by default, being limited to flat 16x16 arrays. While the developers note in their paper that different configurations could be implemented, this would require changes to the OpenMPD code. As such, this does not constitute native support for arbitrary geometries. Additionally, only the piston model propagator is supplied, and thus the more complex scenarios that BEM unlocks [HapticSoftness, HighSpeedBEM] are not possible with OpenMPD as provided. Similarly, only the projective solvers (i.e., GS-PAT, naive and IB) are provided, meaning any scenario that requires a different solver cannot be used.
2.2.3 Levitate
@@
Levitate is a python-based acoustic holography library [Levitate], which allows modelling propagation from transducer arrays to points in space based on the piston model. This library supports an explicit-gradient descent solver where the gradient must be computed by the user and does not integrate any of the more common projective methods or allow for implicit-gradient methods, which may limit its application as these are the dominant methods used across various domains [HapticsMPSTM, LeviPrint, Articulev]. Levitate can compute all of the analysis metrics discussed but has no implementation for hardware support, indicating that it is strictly a software only library. Finally, it also lacks BEM support, giving it similar disadvantages discussed above. This could be a useful software package and maybe one of the most comparable to AcousTools, however, the lack of support for critical sections of the modern holography pipeline, such as many of the most common solvers, would limit its utility.
2.2.4 Other Software
@@
There are various other pieces of acoustic software, which mostly aim to simulate acoustic wave propagation and lack most of the features that we have identified that a developer would need. For example, K-wave is a MATLAB package that enables simulation of free-field acoustics and definition of arbitrary arrays. However, it only supports the most simple geometric solvers, does not have any scattering support and cannot compute any of the more complex analysis metrics[Kwave]. Similar issues are found in other software implementations, such as Ultrasim [ultrasim] and FOCUS [FOCUS], which all lack the versatility that would be needed to give a developer the full range of options. Similarly, there are packages aim at soling BEM problems for example, OpenBEM[OpenBEM], but do not cover any other stage in the stack
@@
Because of the aforementioned reasons, these packages are not considered in the remaining of this work as they do not provide the user with enough of the full-stack to allow for flexible development. For a (non-exhastive) list of acoustic simulation software, which share many of these issues, see Ref.[kwave-software-list].
@@
Alternatively, some software acts as a single application for acoustic development instead of a library for development. One such pipeline is ArticuLev[Articulev], which allows developers to use acoustically levitated props to create display content. ArticuLev provides support for the creation of traps and props and it’s methods are agnostic to the underlying algorithms that are used to generate these traps but only a multipoint version of the naive algorithm is actually provided and in a similar way, they are agnostic to the specific array but assume a 16x16 array. This is not strictly comparable to AcousTools as this is a platform built for a specific application instead of a general purpose software library but should be mentioned that for developers with specific goals, specific software may exist [Articulev].
3 The Acoustic ‘Full-Stack’
@@ 3.0.1 Setup
The bottom of the stack is the setup phase, defining the transducer array’s shape and parameters as well as the locations in space where acoustic waves are focused [LeviPrint, AcousticElements].
@@
The transducer arrays used across acoustic applications vary from flat square boards [OpenMPD] to rectangle [AUTD3] and curved surfaces [TinyLev]. Even more complex geometries have also been investigated in theory, such as spherical or triangular arrays, but these are rarely used in practice due to the added complexity [LeviPrint]. Because of the variety of different options, the first decision that a developer must make is which of these arrays to use, and the software they use should accommodate this.
@@
After defining the array, still as part of setup, the developer needs to conceptually define where in space the acoustic waves should be focused. The area a given transducer array can focus will vary based on its geometry. For example, a single-axis device [TinyLev] has significantly less freedom in focusing locations compared to a larger 3D array [OpenMPD, AUTD3]. Therefore, the transducer geometry needs to be accounted for when defining where the focal points can be created.
@@
Finally, the various acoustic parameters and constants need to be defined, such as the frequency of sound (most often 40kHz[OpenMPD, Ultraino, AUTD3, GSPAT, VolDisplaySingleParticle, HapticsMPSTM, LeviPrint]), the reference pressure of transducers at 1m [GSPAT], the density of the medium and levitated objects, and the speed of sound in that medium [BruusAcoustofluidics]. These are often not choices a developer would actively make, but rather physical properties of the system they have selected, for example the frequency of waves used are dependant on the resonant frequency of the transducers. However, these parameters should be acknowledged, as earlier design decisions will influence them.
3.0.2 Propagators
@@
Once the transducers and points have been defined, the next step is to couple them together by modelling how acoustic waves travel from transducers to points [GSPAT, Diff-PAT, HapticsMPSTM, HighSpeedBEM, ToBEMorNot, HapticSoftness]. There are two main scenarios that can be used, free-field with no scattering objects and propagation with obstacles acting as scatterers. The free-field case, using the piston model as described in (1), is fairly easy to compute and can be calculated at very high speeds for real-time dynamic holography [GSPAT]. It is versatile and, in most free-field cases, the scattering effects can be neglected as they do not impact performance[GSPAT, LeviPrint].
@@
However, in some cases, scattering effects are too significant to ignore [HighSpeedBEM, ToBEMorNot]. Thus, the developer may choose to use a propagator that models this scattering. Given the model of the scatterer as a mesh, BEM can be used to predict how the sound scatterers off the boundary of the object [HighSpeedBEM] as seen in (3). Accounting for them can bring improvements in a multitude of applications [HapticSoftness, ToBEMorNot, HighSpeedBEM] but come at a significant computational cost [ToBEMorNot].
@@ 3.0.3 Solvers
Once the setup and propagator have been defined, acoustic solvers compute the transducer activations that cause transducers to focus sound at the chosen points [VolDisplaySingleParticle, GSPAT, ToBEMorNot]. Supporting a number of solvers enables a developer to have different options, but determining the best solver requires careful analysis of the problem the developer is trying to solve. For an application that requires fast computation speeds, an algorithm such as GS-PAT[GSPAT] may be best. On the other hand, WGS [TGS] gives high pressure and low variance across multiple points. For more complex scenarios, a implicit-gradient descent optimiser can solve more complex objectives [Diff-PAT]. It is important for a developer to be able to define custom objective functions for their own context and the developer must weight which is best, potentially comparing multiple options before a single best choice is decided upon. This emphasises the desirability for simple swapping between solvers, to aid in this development process.
3.0.4 Analysis
@@
At this point, a developer has defined their setup, propagators, and solvers to generate acoustic focal points at the desired locations. After this, they may need to quantify certain properties of the sound field (e.g., pressure, Gor’kov potential, force, stiffness), either for validation or further optimisation [AcousticElements, AcoustoFab, HighSpeedBEM, LeviPrint]. Importantly, this step may overlap with the Solvers step. For example, using a custom objective defined to maximise the pressure at a point may require elements from both the previous and current layer. However, this layer can also sit by itself, for example, quantifying the force on a particle to validate if a particle can be correctly levitated at the desired location [Ultraino]. The quantities discussed here, which help to quantify an acoustic field, are pressure as well as Gor’kov potential, force, and stiffness as described in (8), (11), and (12) respectively.
@@
Which one of these metrics is the most useful will depend on the use case a specific developer might have. If they are developing a haptic interface, the pressure at a point might be the only thing that they need to consider [HapticsMPSTM] while for levitation the Gor’kov potential, force, and stiffness might become more appropriate [HighSpeedBEM, AcousticElements]. Additionally, the square of pressure correlates very well with stiffness, and as such, the simpler metric can be used as a proxy for the more complex one in practice [GSPAT]. As discussed above, having a diverse set of tools means that the correct tool can be used at the right moment.
3.0.5 Hardware
@@
The final step before an acoustic holography application can be deployed is to control the physical transducers in the real world. This involves sending synchronised phase and amplitude values from a computer to the device, physically realising the until-this-point simulated fields. This presents difficulties as for the very high update rates that acoustic holography often demands, up to 10kHz for levitation, a high degree of synchronization is needed. Various platforms exist to address this problem. Such platforms include OpenMPD[OpenMPD], which can run at 10kHz; AUTD3 [AUTD3], which is able to run at 1kHz; and Ultraino[Ultraino], which runs at 25kHz. These different platforms also have different ways to communicate between the transducer arrays and the computer, which would need to be accommodated by any software used.
@@
This section is more difficult for software packages to be general purpose, as each different hardware setup has a different communication method and structure for the messages to be sent [OpenMPD, AUTD3]. Thus, many software packages only function for one or a small number of hardware devices but having at least one supported hardware device enables the simulated fields to be physicalised in real world.
4 AcousTools
Figure 3: Examples of using different meshes as the definition for transducer arrays creating a focal point in the origin of the working area (high pressure region at the centre of each subfigure). For a given mesh, a transducer is placed at each cell centre with the normal of the mesh defining the transducer normal, with normals pointing into the working volume inside the mesh. a) flat array similar to a common setup used in practice [AcousticElements], b) Sphere c) Teapot. For each array, the mesh is scaled so that the limits of the mesh in the x-axis is $\pm$11.8cm, centred at the origin (except the flat array which is translated down in order to give a working volume above it) and the normals point towards the origin.
@@ 4 AcousTools
As discussed above, the existing software implementations for acoustic holography are limited in their functionality, often not supporting a full range of solvers or propagators. To address this, we present AcousTools; a python library built to meet the full-stack of acoustic holography. AcousTools is open source and can be found at our GitHub repository[AcousTools].
@@
Python was chosen to implement AcousTools because of its reputation for accessibility, especially for novice users [pythonEase]. It is one of the most widely used programming languages [PythonUse] and among the fastest for beginners to learn. Whereas C++, commonly used in acoustic holography, is considered one of the most difficult [PythonLearning]. This popularity, combined with Python’s ease of adoption and of use, suggests that adoption and extension of AcousTools would be similarly straightforward.
@@
Additionally, AcousTools is based on PyTorch[PyTorch] due to its parallelised Tensor implementations, GPU support, and automatic differentiation system. These together make PyTorch an ideal base for AcousTools and provide the secondary benefit that potential future machine leaning applications for acoustic holography could be easily implemented, as PyTorch is widely used in this domain.
@@
Here we discuss AcousTools’ ability to meet every step of every stage of the acoustic full-stack discussed above. Table 1 summarises both the existing software’s performance and AcousTools’ features, illustrating that AcousTools meet every stage while the others all lack in some area or another. There is also a lack of standardisation across acoustic holography research [OpenMPD, AUTD3, Ultraino], and we believe that AcousTools with its general purpose nature could fill this gap for researchers across all domains.
4.1 Setup